# Identification of Dynamic Nonlinear Panel Models

### v2

@@ 3.1 Model and Identification
Under Assumption 1, the identified set $\Theta_{I,2}$ for $\theta_{0}$ is the set of parameters $\theta$ that satisfy the following conditions:
$1-\sum_{j=1}^{J}\Pr(Y_{is}=y_{j},b_{j}-z_{s}^{\prime}\beta-X_{is}^{\prime}\gamma\geq a\mid z_{st})\geq\sum_{j=1}^{J}\Pr(Y_{it}=y_{j},b_{j+1}-z_{t}^{\prime}\beta-X_{it}^{\prime}\gamma\leq a\mid z_{st}),$ (9)
for any $a\in\mathcal{R}$, any $s,t\leq T$, and any $z_{st}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{z}}$.
Proposition 4 characterizes the identified set for $\theta_{0}$ for the general ordered response model with potential endogeneity. This result allows for both static and dynamic models, as discussed in Section 3.2. In comparison to the binary choice model, Proposition 4 aggregates information from all $J$ ordered responses across different time periods to identify $\theta_{0}$. When there are only two choices ($J=2$), this result simplifies to the identifying condition (7) in Proposition 1, with $y_{1}=0,y_{2}=1$ and changing $a$ to $-a$. Again, the result in Proposition 4 accommodates both discrete and continuous endogenous covariates. When the endogenous $X_{it}$ only takes a finite number of values, such as lagged dependent variable, we can reduce the number of restrictions in Proposition 4 to a finite number, as shown in the next corollary.
@@
Under Assumption 1 and $X_{it}$ only takes finite number of values $X_{it}\in\{a_{1},...,a_{K}\}$, the identified set $\Theta_{I,2}$ is the set of parameters $\theta$ that satisfy condition (9) for any $a\in\mathcal{Q}_{2}:=\{b_{j}-z_{s}^{\prime}\beta-a_{1}^{\prime}\gamma,...,b_{j}-z_{s}^{\prime}\beta-a_{K}^{\prime}\gamma,b_{j}-z_{t}^{\prime}\beta-a_{1}^{\prime}\gamma,...,b_{j}-z_{t}^{\prime}\beta-a_{K}^{\prime}\gamma\}_{j=2}^{J}$, any $s,t\leq T$, and any $z_{st}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{z}}$.
Similar to Corollary 1, Corollary 2 reduces the number of conditional moment inequalities to $2K\times(J-1)$ numbers for any pair of two periods. For the dynamic model with one lagged dependent variable $X_{it}=Y_{i,t-1}\in\{y_{1},...,y_{J}\}$, then there will be $2J(J-1)$ number of restrictions.
3.2 Applications
In this section, we apply the results in Proposition 4 to explore two panel ordered choice model. The first one examines a static model where all covariates are exogenous, while the other studies a dynamic model with one lagged dependent variable. To the best of our knowledge, these two scenarios have not been explored in the literature within the framework of the stationarity assumption.
Static model: we consider a static model where Assumption 1 holds conditional on all covariates $W_{ist}$. The identifying restriction in Proposition 4 is given as
$1-\sum_{j=1}^{J}\Pr(Y_{is}=y_{j}\mid w_{st})\mathbbm{1}\{b_{j}-w_{s}^{\prime}\theta_{0}\geq a\}\\ \geq\sum_{j=1}^{J}\Pr(Y_{it}=y_{j}\mid w_{st})\mathbbm{1}\{b_{j+1}-w_{t}^{\prime}\theta_{0}\leq a\}.$
We can further simplify the above condition by getting rid of the parameter $a$ and transform it into finite number of conditional moment inequalities. The following lemma presents the identified set for $\Theta_{I,2}$ for this static model.
Lemma 1.
Assume that $\epsilon_{is}\mid(W_{ist},\alpha_{i})\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid(W_{ist},\alpha_{i})$, the identified set $\Theta_{I,2}$ for $\theta_{0}$ is the set of parameters $\theta$ that satisfy the following conditions:
$b_{j_{1}+1}-w_{s}^{\prime}\theta\geq b_{j_{2}+1}-w_{t}^{\prime}\theta\Longrightarrow\sum_{j=1}^{j_{1}}\Pr(Y_{is}=y_{j}\mid w_{st})\geq\sum_{j=1}^{j_{2}}\Pr(Y_{it}=y_{j}\mid w_{st}),$
for any $11\leq j_{1},j_{2}\leq J-$ 1, $s,t\leq T$, and $w_{st}\in\mathcal{R}^{d}\times\mathcal{R}^{d}$.
In the static model, variation in conditional choice probability only comes from changes in the covariate $W_{it}$. Thus, we can directly establish the relationship between the variation in choice probability and the changes in covariates over time, eliminating the parameter $a$. Different from the binary choice model, the results in Lemma 1 also exploit intertemporal variation in the sum of multiple choices rather than investigating a single choice. Moreover, we can utilize variations in the sum of different choices across various periods to identify $\theta_{0}$.
Besides the static model, the results in Proposition 4 can also be applied to study a dynamic ordered response model, where people’s choice at the current period ($t$) can depend on their choice in the last period ($t-1$).
Dynamic model: consider the following dynamic ordered response model with one lagged dependent variable:
@@ 3.2 Applications
In this example, the endogenous covariate is the lagged dependent variable $X_{it}=Y_{it-1}\in\{y_{1},...,y_{J}\}$. The identifying restriction in Proposition 4 is presented as follows:
$1-\sum_{j=1}^{J}\Pr\left(Y_{is}=y_{j},b_{j}-z_{s}^{\prime}\beta-Y_{is-1}\gamma\geq a\mid z_{st}\right)\\ \geq\sum_{j=1}^{J}\Pr\left(Y_{it}=y_{j},b_{j+1}-z_{t}^{\prime}\beta-Y_{it-1}\gamma\leq a\mid z_{st}\right),$
for any $a\in\{b_{j}-z_{s}^{\prime}\beta-y_{1}\gamma,...,b_{j}-z_{s}^{\prime}\beta-y_{J}\gamma,b_{j}-z_{t}^{\prime}\beta-y_{1}\gamma,...,b_{j}-z_{t}^{\prime}\beta-y_{J}\gamma\}_{j=2}^{J}$.
For the dynamic model, the identified set $\Theta_{I,2}$ is characterized by $2J(J-1)$ number of conditional inequalities for any pair of two periods $(s,t)$. This approach can also allow for dynamic models with more than one lagged dependent variable, such as $X_{it}=(Y_{i,t-1},Y_{i,t-2})$. Furthermore, it can accommodate dynamic models that allow for heterogeneous effects from different choices in the last period, as given below:
$Y_{it}^{*}=Z_{it}^{\prime}\beta_{0}+\sum_{j=2}^{J}\mathbf{\mathbbm{1}}\{Y_{it-1}=j\}\gamma_{0,j}+\alpha_{i}+\epsilon_{it}.$
3.3 Point Identification
This section provides sufficient conditions to achieve point identification for $\beta_{0}$ (up to scale) and the sign of $\gamma_{0}$. Similar to Section 2.3, we still focus on the scenario where the endogenous covariate $X_{it}$ only takes finite number of values $X_{it}\in\mathcal{X}=\{a_{1},a_{2},...,a_{K}\}$, and we illustrate the concept using two periods ($T=2$).
To identify $\beta_{0}$, we define the following two sets:
@@ 3.3 Point Identification
$1-\sum_{j=k}^{J}\Pr(Y_{i1}=y_{j},X_{i1}=x\mid z_{12})<\sum_{j=1}^{k-1}\Pr(Y_{i2}=y_{j},X_{i2}=x\mid z_{12})\Big{\}};$
[…]
$1-\sum_{j=1}^{k-1}\Pr(Y_{i1}=y_{j},X_{i1}=x\mid z_{12})<\sum_{j=k}^{J}\Pr(Y_{i2}=y_{j},X_{i2}=x\mid z_{12})\Big{\}}.$
Let $\mathcal{Z}_{order}:=\mathcal{Z}_{1,order}\cup\mathcal{Z}_{2,order}$. Let $\Delta Z_{i}=Z_{i2}-Z_{i1}$ and $\Delta\mathcal{Z}_{order}$ be defined as
$\Delta\mathcal{Z}_{order}:=\Big{\{}\Delta z:=z_{2}-z_{1}\mid(z_{1},z_{2})\in\mathcal{Z}_{order}\Big{\}}.$
Assumption 3 (Support Condition).
(1) $\Delta\mathcal{Z}_{order}$ is not contained in any proper linear subspace of $\mathcal{R}^{d_{z}}$; (2) for any $\Delta z\in\Delta\mathcal{Z}_{order}$, there exists at least one least one element $\Delta z^{j^{*}}$ with $\beta_{0}^{j*}\neq 0$, and the conditional support of $\Delta z^{j^{*}}$ is $\mathcal{R}$ given $\Delta z\setminus\Delta z^{j^{*}}$, where $\Delta z\setminus\Delta z^{j^{*}}$ denote the remaining components of $\Delta z$ besides $\Delta z^{j^{*}}$.
@@
Under Assumptions 1 &\&& 3, the parameter $\beta_{0}$ is point identified up to scale.
Similar to Proposition 2, the strategy is to first identify the sign of $\Delta z^{\prime}\beta_{0}$ from certain variation in the choice probability, as defined in the two sets $(\mathcal{Z}_{1,order},\mathcal{Z}_{2,order})$. Given the sign of $\Delta z^{\prime}\beta_{0}$ is identified, then the parameter $\beta_{0}$ is point identified up to scale under the conventional large support condition on $\Delta z$. For the static model (with $\gamma_{0}=0$), all coefficients $\beta_{0}$ are point identified under the support condition.
To identify the sign of $\gamma_{0}$, we define the following two sets: for $j\leq d_{x}$,
[…]
$1-\sum_{j=k}^{J}\Pr(Y_{i1}=y_{j},X_{i1}=x_{1}\mid z_{12})<\sum_{j=1}^{k-1}\Pr(Y_{i2}=y_{j},X_{i2}=x_{2}\mid z_{12})\Big{\}};$
[…]
$1-\sum_{j=1}^{k-1}\Pr(Y_{i1}=y_{j},X_{i1}=x_{1}\mid z_{12})<\sum_{j=k}^{J}\Pr(Y_{i2}=y_{j},X_{i2}=x_{2}\mid z_{12})\Big{\}}.$
Proposition 6.
Under Assumptions 1, and for any $1\leq j\leq d_{x}$, either $\mathcal{Z}^{j}_{3,order}\cap\mathcal{Z}_{2,order}\neq\emptyset$ or $\mathcal{Z}^{j}_{4,order}\cap\mathcal{Z}_{1,order}\neq\emptyset$, then the sign of $\gamma_{0}$ is identified.
For the dynamic model with one lagged dependent variable $X_{it}=Y_{i,t-1}$, then the joint choice probabilities defined in the two sets $\mathcal{Z}^{j}_{3,order},\mathcal{Z}^{j}_{4,order}$ can be simplified as taking different values for the lagged dependent variable: $Y_{i0}=y_{j_{1}},Y_{i1}=y_{j_{2}}$ with $y_{j_{1}}\neq y_{j_{2}}$. The identifying restriction in Proposition 4, in general, can bound the value of $(y_{j}-y_{k})\gamma_{0}$ using the variation in $Z_{it}$. Proposition 6 ensures that its bound excludes zero for some $(y_{j},y_{k})$, thereby identifying the sign of $\gamma_{0}$. When there are more than two periods ($T>2$), then point identification of $\beta_{0}$ and the sign of $\gamma_{0}$ is achieved when there exists two periods such that the assumptions are satisfied.
4 Multinomial Choice Model
4.1 Model and Identification
This section applies our key identification strategy to panel multinomial choice model with endogeneity. Specifically, consider a set of choice alternatives $\mathcal{J}=\{0,1,...,J\}$. Let $u_{ijt}$ denote the utility for individual $i$ for choice $j$ at time $t$, which depends on three components: observed covariate $W_{ijt}=(Z_{ijt}^{\prime},X_{ijt}^{\prime})^{\prime}$, unobserved fixed effects $\alpha_{ij}$, and unobserved time-varying preference shock $\epsilon_{ijt}$. The utility of choice 0 (outside option) is normalized to zero: $u_{i0t}=0$. Let $Y_{it}\in\mathcal{J}$ denote individual $i$’s choice at time $t$. We study the following panel multinomial choice model:
$u_{ijt}$ $=Z_{ijt}^{\prime}\beta_{0}+X_{ijt}^{\prime}\gamma_{0}+\alpha_{ij}+\epsilon_{ijt},$
$Y_{it}$ $=\arg\max_{j\in\mathcal{J}}u_{ijt},$
where $X_{ijt}$ denotes the potentially endogenous covariate, such as the lagged dependent variable or endogenously determined prices, and $Z_{ijt}$ denotes the exogenous covariates that satisfies the partial stationarity condition (Assumption 1)
$\epsilon_{is}\mid Z_{is},Z_{it},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid Z_{is},Z_{it},\alpha_{i}\quad\text{for any}\ s,t\leq T.$
with $Z_{it}=\{Z_{ijt}\}_{j\in\mathcal{J}}$,$\alpha_{i}=\{\alpha_{ij}\}_{j\in\mathcal{J}}$ and $\epsilon_{it}=\{\epsilon_{ijt}\}_{j\in\mathcal{J}}$ defined to collect terms across all $J$ choice alternatives.
The identification of $\theta_{0}$ has been studied under the standard full stationarity condition given all covariates $W_{ist}$ in different models, including Pakes and Porter (2022), Shi, Shum, and Song (2018), Gao and Li (2020), and Wang (2022). The main focus of this paper is to derive identification for $\theta_{0}$, while allowing for the endogenous covariate $X_{it}$. As discussed in Section 4.2, our approach also accommodates the standard stationarity assumption, and the results are consistent with those in Pakes and Porter (2022).
Let $v_{ijt}=\alpha_{ij}+\epsilon_{ijt}$ and $v_{it}=\{v_{ijt}\}_{j\in\mathcal{J}}$. The identification strategy is still to establish bounds on the distribution of $v_{it}\mid Z_{ist}$ and derive identifying restrictions on $\theta_{0}$ based on partial stationarity. Here, the distinction is that the error term $v_{it}$ is multi-dimensional instead of one-dimensional in binary or ordered choice models. Moreover, consumers’ choice involves the comparison between different choices.
The observed choice probability of selecting $j$ at period $t$ is given as
@@ 4.1 Model and Identification
$\Pr(Y_{it}=j\mid w_{st})=\Pr\left(w_{jt}^{\prime}\theta_{0}+v_{ijt}\geq w_{kt}^{\prime}\theta_{0}+v_{ikt}\quad\forall k\in\mathcal{J}\mid w_{st}\right).$
The above probability restricts the distribution $\Pr(v_{ikt}-v_{ijt}\leq(w_{jt}-w_{kt})^{\prime}\theta_{0}\ \forall k\neq j\mid w_{st})$. From observed choice probability, we can derive bounds for the distribution $\Pr(v_{ikt}-v_{ijt}\leq a_{jk}\ \forall k\neq j\mid z_{st})$ as a function of $\theta_{0}$ and observed variable $(X_{it},Z_{it},Y_{it})$. Therefore, the identifying condition for $\theta_{0}$ is characterized by the restriction that the bounds over different periods must intersect. More generally, we observe the conditional probability $\Pr(Y_{it}\in K\mid w_{st})$ for any set $K\subset\mathcal{J}$ and can bound the probability $\Pr(v_{ikt}-v_{ijt}\leq a_{jk}\ \exists j\in K,\forall k\in K^{c}\mid z_{st})$, where $K^{c}:=\mathcal{J}\setminus K$. In the case where $K$ is a singleton ($K=\{j\}$), it then becomes the conditional probability of selecting a single choice $j$. The following proposition characterizes the identification results for $\theta_{0}$.
Proposition 7.
Under Assumption 1, an identified set $\Theta_{I,3}$ for $\theta_{0}$ is the set of parameters $\theta=(\beta,\gamma)$ that satisfy the following condition:
$1-\Pr(Y_{is}\in K^{c},(z_{js}-z_{ks})^{\prime}\beta+(X_{ijs}-X_{iks})^{\prime}\gamma\geq a_{jk}\ \forall j\in K,k\in K^{c}\mid z_{st})\\ \geq\Pr(Y_{it}\in K,(z_{jt}-z_{kt})^{\prime}\beta+(X_{ijt}-X_{ikt})^{\prime}\gamma\leq a_{jk}\ \forall j\in K,k\in K^{c}\mid z_{st}),$ (10)
for any subset $K\subset\mathcal{J}$, any $a_{jk}\in\mathcal{R}$ with $j\in K$ and $k\in K^{c}$, any $s,t\leq T$, and any $z_{st}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{z}}$.
Corollary 3.
Under Assumption 1, and when $X_{ijt}-X_{ikt}$ only takes finite number of values $X_{ijt}-X_{ikt}\in\mathcal{X}_{jk}:=\{x_{jk,1},...,x_{jk,M}\}$, then the identified set $\Theta_{I,3}$ is the set of parameters $\theta$ that satisfy condition (10) for any subset $K\subset\mathcal{J}$, any $a_{jk}\in\{(z_{js}-z_{ks})^{\prime}\beta+x_{jk}^{\prime}\gamma,(z_{jt}-z_{kt})^{\prime}\beta+x_{jk}^{\prime}\gamma\}_{x_{jk}\in\mathcal{X}_{jk}}$ with $j\in K$ and $k\in K^{c}$, any $s,t\leq T$, and any $z_{st}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{z}}$.
Proposition 7 provides general identification results for multinomial choice models with discrete or continuous endogenous covariates. Corollary 3 further reduces the number of restrictions in Proposition 7 to a finite number when the difference in the endogenous covariate between choices, $X_{ijt}-X_{ikt}$, only takes finite number of values. The results accommodate both static models and dynamic models with any number of lagged dependent variables. Point identification can be achieved under conditions similar to those presented in Section 2.3 and 3.3, so further analysis is omitted here.
4.2 Applications
We examine the two applications of Proposition 7 for the panel multinomial choice model. The first one studies the static model, where all covariates $W_{it}$ are exogenous ($\epsilon_{is}\mid W_{ist},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid W_{ist},\alpha_{i}$), which is the same setting as Pakes and Porter (2022). In this scenario, we show that Proposition 7 boils down to the identifying results in Pakes and Porter (2022). Another one explores the dynamic multinomial choice model, incorporating one lagged dependent variable $Y_{i,t-1}$.
@@ 4.2 Applications
Static model: Pakes and Porter (2022) characterizes the sharp identified set for $\theta_{0}$ under the stationarity assumption given all covariates: $\epsilon_{is}\mid W_{is},W_{it},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid W_{is},W_{it},\alpha_{i}$. Now, we show that the identifying restriction in (10) is consistent with the results in Pakes and Porter (2022).
Under full stationarity (given all covariates $W_{it}$), condition (10) becomes
$1-\Pr(Y_{is}\in K^{c},(w_{js}-w_{ks})^{\prime}\theta_{0}\geq a_{jk}\ \forall j\in K,k\in K^{c}\mid w_{st})\\ \geq\Pr(Y_{it}\in K,(w_{jt}-w_{kt})^{\prime}\theta_{0}\leq a_{jk}\ \forall j\in K,k\in K^{c}\mid w_{st}).$
@@
The above equation is only informative when $a_{jk}\in[(w_{jt}-w_{kt})^{\prime}\theta_{0},(w_{js}-w_{ks})^{\prime}\theta_{0}]$ for any $j\in K,k\in K^{c}$; otherwise either the upper bound becomes one or the lower bound becomes zero in the above restriction. When $a\in[(w_{jt}-w_{kt})^{\prime}\theta_{0},(w_{js}-w_{ks})^{\prime}\theta_{0}]$, the above restriction is equivalent to the following condition: for any $K\subset\mathcal{J}$,
@@
$\text{If }(w_{js}-w_{ks})^{\prime}\theta_{0}\geq(w_{jt}-w_{kt})^{\prime}\theta_{0}\ \ \forall j\in K,k\in K^{c}\\ \Longrightarrow 1-\Pr(Y_{is}\in K^{c}\mid w_{st})=\Pr(Y_{is}\in K\mid w_{st})\geq\Pr(Y_{it}\in K\mid w_{st}),$
which aligns with the identification result in Pakes and Porter (2022) (Proposition 1, P. 12).
Dynamic model: Proposition 7 is also applicable to the following dynamic panel multinomial choice model:
@@
The above model allows individuals’ current utility at time $t$ to depend on their choices at the last period $t-1$. In this model, the endogenous variable $X_{ijt}$ is the lagged dependent variable $Y_{i,t-1}$, and the difference of the endogenous covariate between choices only takes three values: $X_{ijt}-X_{ikt}\in\mathcal{X}_{jk}:=\{1,-1,0\}$. Then the identified set for $\Theta_{I,3}$ is characterized by the condition in Corollary 2 with $a_{jk}\in\{(z_{js}-z_{ks})^{\prime}\beta+\gamma,(z_{js}-z_{ks})^{\prime}\beta-\gamma,(z_{js}-z_{ks})^{\prime}\beta,(z_{jt}-z_{kt})^{\prime}\beta+\gamma,(z_{jt}-z_{kt})^{\prime}\beta-\gamma,(z_{jt}-z_{kt})^{\prime}\beta\}$.
5 Censored Dependent Variable Model
The previous sections primarily investigate discrete choice models, while our approach also applies to models with continuous dependent variables, including those with censored or interval outcomes. To illustrate, we focus on censored outcomes models below.
Consider the following panel models with censored outcomes:

### v3

@@ 3.2 Ordered Response Model
Under Assumptions 1, $\theta_{0}\in\Theta_{I,order}$, where the identified set $\Theta_{I,order}$ consists of all $\theta=\left(\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right)^{{}^{\prime}}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{x}}$ such that
$\max_{t=1,...,T}\sum_{j=1}^{J}\mathbb{P}\left(Y_{it}=y_{j},b_{j+1}-z_{t}^{\prime}\beta-X_{it}^{\prime}\gamma\leq c\mid z\right)\leq\\ 1-\max_{s=1,...,T}\sum_{j=1}^{J}\mathbb{P}\left(Y_{is}=y_{j},b_{j}-z_{s}^{\prime}\beta-X_{is}^{\prime}\gamma\geq c\mid z\right),$ (22)
for any $c\in\mathcal{R}$ and any realization $z=(z_{1},...,z_{T})$ in the support of $Z_{i}$.
Proposition 7 characterizes the identified set of $\theta_{0}$ for the general ordered response model with endogeneity. In contrast to the binary choice model discussed in Section 2, Proposition 7 exploits information from all possible choices across different time periods to identify $\theta_{0}$. This result accommodates both static and dynamic models, and we derive the simplified identifying results for these two types of models as follows.
Static model: consider that there is no endogeneity and the full stationarity assumption holds conditional on all covariates $W_{i}$:
$\epsilon_{is}\mid W_{i},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid W_{i},\alpha_{i}.$
The identifying restriction in Proposition 7 is given as
$\max_{t=1,...,T}\sum_{j=1}^{J}\mathbb{P}\left(Y_{it}=y_{j},\ b_{j+1}-w_{t}^{\prime}\theta_{0}\leq c\mid w\right)\leq 1-\max_{s=1,...,T}\sum_{j=1}^{J}\mathbb{P}(Y_{is}=y_{j},\ b_{j}-w_{s}^{\prime}\theta_{0}\geq c\mid w).$
The above condition is informative only if there exists $j_{1},j_{2}$ such that $b_{j_{2}+1}-w_{t}^{\prime}\theta_{0}\leq c\leq b_{j_{1}}-w_{s}^{\prime}\theta_{0}$, leading to
$\max_{t=1,...,T}\sum_{j=1}^{j_{2}}\mathbb{P}\left(Y_{it}=y_{j}\mid w_{i}\right)$ $\leq 1-\max_{s=1,...,T}\sum_{j=j_{1}}^{J}\mathbb{P}(Y_{is}=y_{j}\mid w)$
$=\max_{s=1,...,T}\sum_{j=1}^{j_{1}-1}\mathbb{P}(Y_{is}=y_{j}\mid w).$
The following proposition presents the identification results for the static ordered choice model by changing $j_{1}-1$ with $j_{1}$.
@@
Assume that $\epsilon_{is}\mid(W_{i},\alpha_{i})\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid(W_{i},\alpha_{i})$, then $\Theta_{I,order}$ consists of all $\theta=\left(\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right)^{{}^{\prime}}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{x}}$ such that
$b_{j_{1}+1}-w_{s}^{\prime}\theta\geq b_{j_{2}+1}-w_{t}^{\prime}\theta\Longrightarrow\max_{s=1,...,T}\sum_{j=1}^{j_{1}}\mathbb{P}(Y_{is}=y_{j}\mid w)\geq\max_{t=1,...,T}\sum_{j=1}^{j_{2}}\mathbb{P}(Y_{it}=y_{j}\mid w),$
for any $11\leq j_{1},j_{2}\leq J-$ 1, and any realization $w=(w_{1},...,w_{T})$ in the support of $W_{i}$.
The results in Corollary 1 are analogous to the maximum-score type result in Manski (1987), with the distinction being that we can exploit variations in the sum of multiple choices rather than investigating a single choice to identify $\theta_{0}$. Furthermore, with multiple choices, we can utilize variations in the sum of different choices across various periods for identification. Besides the static model, Proposition 7 can also accommodate dynamic ordered choice models with lagged dependent variable.
Dynamic model: consider the following dynamic ordered choice model with one lagged dependent variable:
@@
3.3 Multinomial Choice Model
This section applies our key identification strategy to panel multinomial choice model with endogeneity. Specifically, consider a set of unordered choice alternatives $\mathcal{J}=\{0,1,...,J\}$. Let $u_{ijt}$ denote the latent utility for individual $i$ of selecting choice $j$ at time $t$, which depends on the three components: observed covariate $W_{ijt}=(Z_{ijt}^{\prime},X_{ijt}^{\prime})^{\prime}$, unobserved fixed effects $\alpha_{ij}$, and unobserved time-varying preference shock $\epsilon_{ijt}$. Let $Y_{it}\in\mathcal{J}$ denote individual $i$’s choice at time $t$. We study the following panel multinomial choice model:
$u_{ijt}$ $=W_{ijt}^{\prime}\theta_{0}+\alpha_{ij}+\epsilon_{ijt},$
$Y_{it}$ $=\arg\max_{j\in\mathcal{J}}u_{ijt},$
and impose the same partial stationarity assumption:
$\epsilon_{is}\mid Z_{i},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid Z_{i},\alpha_{i}\quad\text{for any}\ s,t\leq T.$
with $Z_{it}:=\{Z_{ijt}\}_{j\in\mathcal{J}}$,$\alpha_{i}:=\{\alpha_{ij}\}_{j\in\mathcal{J}}$ and $\epsilon_{it}:=\{\epsilon_{ijt}\}_{j\in\mathcal{J}}$ defined to collect terms across all $J$ choice alternatives, and $Z_{i}:=(Z_{i1},Z_{i2},...,Z_{iT})$.
Although the $J$ choices are unordered and not directly comparable, we look at the indicator variable $Y_{it}^{j}:=\mathbf{\mathbbm{1}}\{Y_{it}=j\}$ of choosing option $j$, which maintains a similar monotone structure with Assumption 3:
$Y_{it}^{j}=1\Longleftrightarrow W_{ijt}^{\prime}\theta_{0}+\alpha_{ij}+\epsilon_{ijt}\geq W_{ikt}^{\prime}\theta_{0}+\alpha_{ik}+\epsilon_{ikt},\ \forall k\in\mathcal{J},$
and the new variable $Y_{it}^{j}$ is monotone in $W_{ijt}^{\prime}\theta_{0}-W_{ikt}^{\prime}\theta_{0}$ given $(\alpha_{i},\epsilon_{it})$.
More generally, for any subset $K\subset\mathcal{J}$, the indicator variable $Y_{it}^{K}:=\mathbf{\mathbbm{1}}\{Y_{it}\in K\}$ represents individual $i$’s choice belonging to the subset $K$, given as follows:
$Y_{it}^{K}=1\Longleftrightarrow W_{ijt}^{\prime}\theta_{0}+\alpha_{ij}+\epsilon_{ijt}\geq W_{ikt}^{\prime}\theta_{0}+\alpha_{ik}+\epsilon_{ikt},\ \exists j\in K,\forall k\in\mathcal{J}\setminus K,$
and the variable $Y_{it}^{K}$ is monotone in $W_{ijt}^{\prime}\theta_{0}-W_{ikt}^{\prime}\theta_{0}$ for any $j\in K$ and $k\in\mathcal{J}\setminus K$.
Following Proposition 5, the identification results for panel multinomial choice models are presented in the following proposition.
@@ 3.3 Multinomial Choice Model
Under Assumption 1, $\theta_{0}\in\Theta_{I,mul}$, where the identified set $\Theta_{I,mul}$ consists of all $\theta=\left(\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right)^{{}^{\prime}}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{x}}$ such that
@@
$\max_{t=1,...,T}\mathbb{P}(Y_{it}^{K}=0,(z_{js}-z_{ks})^{\prime}\beta+(X_{ijs}-X_{iks})^{\prime}\gamma\geq c_{jk}\ \forall j\in K,k\in\mathcal{J}\setminus K\mid z)\\ \leq 1-\max_{s=1,...,T}\mathbb{P}\left(Y_{is}^{K}=1,(z_{js}-z_{ks})^{\prime}\beta+(X_{ijs}-X_{iks})^{\prime}\gamma\leq c_{jk}\ \forall j\in K,k\in\mathcal{J}\setminus K\mid z\right),$ (23)
for any subset $K\subset\mathcal{J}$, any $c_{jk}\in\mathcal{R}$, any $j\in K$ and $k\in\mathcal{J}\setminus K$, and any realization $z=(z_{1},...,z_{T})$ in the support of $Z_{i}$.
Proposition 8 provides general identification results for multinomial choice models with endogenous covariates. It accommodates static, dynamic multinomial choice models with any number of lagged dependent variables, as well as other types of endogeneity. For the dynamic model, similar to Proposition 2, the results can be simplified to hold for a finite number of values of $c_{jk}$, which is the collection of values the covariate index $(z_{jt}-z_{kt})^{\prime}\beta+(X_{ijt}-X_{ikt})^{\prime}\gamma$ can take.
Reconciliation with Pakes and Porter (2022)
Next, we show that our results specialize to those in Pakes and Porter (2022), who focuses on the static panel multinomial choice model without any endogeneity.
@@
Formally, Pakes and Porter (2022) characterizes the sharp identified set for $\theta_{0}$ under the full stationarity assumption given all covariates:
$\epsilon_{is}\mid W_{i},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid W_{i},\alpha_{i}.$
Under full stationarity, for any two periods $(s,t)$, our identifying condition in (23) is simplified as
$\mathbb{P}\left(Y_{it}^{K}=0,(w_{jt}-w_{kt})^{\prime}\theta_{0}\geq c_{jk}\ \forall j\in K,k\in\mathcal{J}\setminus K\mid w\right)\\ \leq 1-\mathbb{P}\left(Y_{is}^{K}=1,(w_{js}-w_{ks})^{\prime}\theta_{0}\leq c_{jk}\ \forall j\in K,k\in\mathcal{J}\setminus K\mid w\right)$ (24)
@@
The above equation is only informative when $(w_{js}-w_{ks})^{\prime}\theta_{0}\leq c_{jk}\leq(w_{jt}-w_{kt})^{\prime}\theta_{0}$ for any $j\in K,k\in k\in\mathcal{J}\setminus K$; otherwise either the upper bound becomes one or the lower bound becomes zero so condition (24) holds for any $\theta$. There exists one value $c_{jk}$ satisfying the condition $(w_{js}-w_{ks})^{\prime}\theta_{0}\leq c_{jk}\leq(w_{jt}-w_{kt})^{\prime}\theta_{0}$ is equivalent to $(w_{js}-w_{ks})^{\prime}\theta_{0}\leq(w_{jt}-w_{kt})^{\prime}\theta_{0}$, generating the following inequality: for any $K\subset\mathcal{J}$,
@@
$\text{If }(w_{js}-w_{ks})^{\prime}\theta_{0}\leq(w_{jt}-w_{kt})^{\prime}\theta_{0}\ \ \forall j\in K,k\in\mathcal{J}\setminus K\\ \Longrightarrow\mathbb{P}\left(Y_{it}^{K}=0\mid w\right)\leq 1-\mathbb{P}\left(Y_{is}^{K}=1\mid w\right),$
which becomes the same result in Pakes and Porter (2022) (Proposition 1, P. 12):
$\text{If }(w_{js}-w_{ks})^{\prime}\theta_{0}\leq(w_{jt}-w_{kt})^{\prime}\theta_{0}\ \ \forall j\in K,k\in\mathcal{J}\setminus K\\ \Longrightarrow\mathbb{P}\left(Y_{is}\in K\mid w\right)\leq\mathbb{P}\left(Y_{it}\in K\mid w\right),$
since $Y_{it}^{K}=1$ is equivalent to $Y_{it}\in K$ by the definition.
Beyond the static model, Proposition 8 allows for any type of endogeneity including dynamic multinomial models with lagged dependent variable. For example, consider the following dynamic model:
@@
where individual $i$’s utility at time $t$ can potentially depend on their choices in the previous period $t-1$. In this model, the endogenous variable $X_{ijt}$ is whether option $j$ is chosen in the previous period $\mathbbm{1}\{Y_{i,t-1}=j\}$. Then, the difference in the endogenous covariate between choices only takes three values: $X_{ijt}-X_{ikt}\in\{1,-1,0\}$, and the identified set for $\Theta_{I,mul}$ is characterized by the condition in (23) with $c_{jk}\in\{(z_{jt}-z_{kt})^{\prime}\beta+\gamma,(z_{jt}-z_{kt})^{\prime}\beta-\gamma,(z_{jt}-z_{kt})^{\prime}\beta\}_{t=1}^{T}$.
3.4 Censored Outcome Model
The previous sections primarily investigate discrete choice models, while our approach also applies to models with continuous dependent variables, including those with censored or interval outcomes. To illustrate, we focus on the following panel model with censored outcomes: